Please help me if you can. A 15 feet ladder is leaning against a wall. The base of the ladder is 3 feet from the wall. How high above the ground is the top of the ladder?
Answer by nerdybill(7384)

Please help me if you can. A 15 feet ladder is leaning against a wall. The base of the ladder is 3 feet from the wall. How high above the ground is the top of the ladder?
.
The ladder, wall and ground forms a "right triangle". This allows you to apply Pythagorean theorem.
Let x = height top of ladder reaches
then
x^2 + 3^2 = 15^2
x^2 + 9 = 225
x^2 = 216
x = 14.70 feet